OpenStudy (anonymous):

Solve. p2 – p + 1 = 0

OpenStudy (rulnick):

p^2 - p + 1 = 0 p = 1/2 +/- i sqrt(3)/2

OpenStudy (anonymous):

is that the whole thing

OpenStudy (rulnick):

Well, yes, there are two complex solutions, p = 1/2 + i sqrt(3)/2 and p = 1/2 - i sqrt(3)/2.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

p2 – p + 1 = 0 discriminant = 1 - 4 = -3 -> sqrt (-3) = i * sqrt (3) -> p = [1 +/- isqrt (3) ] / 2